OpenStudy (my4babiesandme):

simplify by factoring 4 square root 96x^4y^6.. please help me understand how to do this

OpenStudy (anonymous):

4 * sqrt(6*16) x^2y^3 = 4*4 sqrt6 x^2y^3 = 16 sqrt6 x^2 y^3 i'm assuming that the sqrt is of (96x^4y^6)

OpenStudy (anonymous):

96 = 2*2*2*2*2*3\[\sqrt[4]{96x ^{4}y ^{6}} = \sqrt[4]{2^{4}*6*x ^{4}*y ^{4}*y ^{2}}\] \[= 2*x*y*\sqrt[4]{6y ^{2}} = 2xy*\sqrt[4]{6y ^{2}}\]